OpenStudy (needhelpp101):

.) Perform the indicated operation without the use of a calculator and express each answer in decimal notation. A. 6e+13 B. 6000 C. 0.00006 D. 60,000 The Problem is in the Comments below

OpenStudy (needhelpp101):

Problem Here :D

OpenStudy (anonymous):

ok, so this is just a manipulation of notations. Think of the equation as such: \[\frac{4.2*(10)*(10)*(10)*(10)*(10)*(10)*(10)*(10)*(10) }{7*(10)*(10)*(10)*(10)*(10)} \] There are a few 10s that cancel one another as you can see, so this fraction reduces to \[\frac{ 4.2*(10)*(10)*(10)*(10) }{ 7 }\] we can rearrange this fraction by multiplying 4.2 by one of the 10s on the numerator\[\frac{ 42*(10)*(10)*(10) }{ 7 }\] This fraction may now be written into an equivalent form\[\frac{ 42 }{7 }*(10)*(10)*(10)\] evaluating this we have\[6*10^3\] which is 6000

OpenStudy (anonymous):

The key is to think of scientific notation in a different manner.